OpenStudy (hannahwaller5):

Will Give Medal!!! The function f(x) = −x2 + 50x − 264 models the profit, in dollars, a yoga studio makes for selling monthly memberships, where x is the number of memberships sold and f(x) is the amount of profit. Part A: Determine the vertex. What does this calculation mean in the context of the problem? Part B: Determine the x-intercepts. What do these values mean in the context of the problem?

OpenStudy (janu16):

do youknow how to find vertex?

OpenStudy (hannahwaller5):

Yes im pretty sure lol

OpenStudy (janu16):

so what is it?

OpenStudy (hannahwaller5):

Well first dont we have to put it in vertex form? Or can you find it without doing that...?

OpenStudy (janu16):

yes y = a(x - h)2 + k

OpenStudy (janu16):

thats the formula

OpenStudy (hannahwaller5):

Right so its in vertex form?

OpenStudy (janu16):

n you have to plug it in the formula

OpenStudy (janu16):

no

OpenStudy (hannahwaller5):

Oh ok so ow do you do that?

OpenStudy (janu16):

you havent learned that?

OpenStudy (hannahwaller5):

no... it wasnt in my lesson at least...

OpenStudy (anonymous):

You can see from the chart that function #2 has a minimum value of (-1,-3). We know this because the x values on either side of -1, -2 and 0, provide us with y values that are greater than -3. Now that we know what the minimum value is for function #2, we need to find the minimum for function #1. We can do this by using the first bit of the quadratic formula (-b/2a). Plugging in function #1 gives us (8/4) or 2. This represents the x value of the vertex. Plugging 2 into function #1 gives us an answer of -7, or a vertex of (2,-7). Now that we know both vertices, we can see that, since -7 is smaller than -3, function #1 has the least minimum value, occurring at (2,-7)

OpenStudy (janu16):

what chart deadpool?

OpenStudy (hannahwaller5):

yea im a little confused... lol

OpenStudy (janu16):

@-Deadpool-

OpenStudy (hannahwaller5):

So i found the vertex and it is (25,361)

OpenStudy (anonymous):

my bad cheerleader I did not mean to say chart, i think i might have read the question wrong lol

OpenStudy (janu16):

thats correct so The vertex represents the maximum possible profit of 361 dollars by selling 25 memberships. .

OpenStudy (janu16):

now part b

OpenStudy (hannahwaller5):

ok

OpenStudy (janu16):

first factor out the negative sign −(x2−50x+264)

OpenStudy (janu16):

so we have (x2−50x+264)

OpenStudy (janu16):

Which two numbers add up to -50 and multiply to 264?

OpenStudy (janu16):

@Hannahwaller5

OpenStudy (hannahwaller5):

ok

OpenStudy (hannahwaller5):

sorry my computer is being weird one sec

OpenStudy (welshfella):

Another way is to complete the square and convert to vertex form- -x^2 + 50 x - 264 = -1(x^2 + 50x) - 264 = -1((x - 25)^2 + 625 - 264 = -1(x _ 25)^2 + 361

OpenStudy (hannahwaller5):

Sorry im not good at finding the common factors and stuff... so im a little confused

OpenStudy (welshfella):

vertex is at (25,361)

OpenStudy (hannahwaller5):

ok so -44 and 6

OpenStudy (janu16):

we already found the vertex

OpenStudy (janu16):

yes

OpenStudy (janu16):

rewrite the expression −(x−44)(x−6) so the x intercets are 44 and 6

OpenStudy (welshfella):

yea i know you found vertex already - i was just showing another way of finding it.

OpenStudy (janu16):

ohok lol

OpenStudy (hannahwaller5):

Oh ok tha makes since thank yall both (:

OpenStudy (janu16):

so those are the x intercepts 6, 44 in the context of this problem, there is positive profit between selling 6 and 44 memberships. Selling less than 6 memberships and greater than 44 memberships there is negative profit or loss.

OpenStudy (janu16):

@Hannahwaller5 thats what it means

OpenStudy (hannahwaller5):

I need to write all this down so i will remeber lol ok thnaks

OpenStudy (janu16):

np